### v3

@@ 2 Background and Definitions
Let $\mathcal{B}(S)$ denote the Borel $\sigma$-algebra on $S$. We consider the logistic Markov process
@@
We can iterate the operator $P$ a positive number $m$ times, which we denote by $P^{m}$: if $X_{0}$ has distribution $\mu$ then $X_{m}$ has distribution $P^{m}\mu$. This leads us to the analog of stability in the stochastic context.
Definition 2.
A sequence of probability measures $P^{n}(n\geq 1)$ on $S$ converges weakly to a probability measure $P$ on $S$ if, for any continuous, bounded function $h:S\mapsto S$,
[…]
In the deterministic case, we know that, if the iterates of $x_{0}$ under a continuous function, $f$, converge to a point, $x$, then $x$ must be a fixed point of $f$, and this result follows from the continuity of $f$. We can get a similar result in the stochastic setting, but we first need to define a notion of continuity. Suppose we have a Markov process on $S$ with transition probabilities given by $p(x,dy)$.
Definition 3.
The transition probability $p(x,dy)$ is weakly continuous or has the Feller property if, for every continuous and bounded function $h:S\mapsto S$, the function
$x\mapsto\int_{S}h(y)p(x,dy)$
is continuous.
Using weak continuity we get the following result ([14], Theorem 11.1 of Chapter 2).
Proposition 1.
Suppose $p(x,dy)$ is a transition probability on $S$ with the Feller property.
  1. 1.
    If there is an $x\in S$, a sequence of integers $1\leq n_{1}<n_{2}<\dots$ and a probability measure $\pi_{x}$ on $S$ such that
    $\frac{1}{n_{k}}\sum_{m=0}^{n_{k}-1}p^{m}(x,dy)\hbox{ converges weakly to }\pi_{x}\hbox{ as }k\rightarrow\infty$
    then $\pi_{x}$ is an invariant measure for $p$.
  2. 2.
    If, for all $x\in S$, there is a sequence of integers $\{n_{i}\}$ such that the limit above exists and is independent of $x$ then $\pi=\pi_{x}$ is the unique invariant probability for $p$.
We can state this proposition in terms of the operator $P$ given in Equation (2). Suppose a process has the Feller property and we have a measure $\mu$ on $S$. If we have weak convergence in terms of the average to a measure, $\pi$:
$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{m=0}^{n-1}P^{m}\mu(A){\rightarrow}\pi(A)$
then $\pi$ is an invariant measure under the operator $P$, i.e. $P\pi(A)=\pi(A)$ for $A\in{\cal B}(S)$. Furthermore, if the condition in Item 2 of Proposition 1 holds, then this invariant measure is unique.
@@
    The Markov process is $\phi$-recurrent if, for each $x\in S$ and $A\in\mathcal{B}(S)$ such that $\phi(A)>0$, we have
@@
    If the Markov process is both $\nu$-irreducible and $\phi$-recurrent, then it is Harris recurrent.
@@
    for some $\phi$-recurrent set $A_{0}$, then the process is said to be positive Harris recurrent.
Note: the set $A_{0}$ is $\phi$-recurrent if $\phi(A_{0})>0$ and it is recurrent, i.e. $P(\eta_{A_{0}}<\infty|X_{0}=x)=1$ $\forall x\in{\mathcal{B}}(S)$.
In other words: a Markov process is $\nu$-irreducible if there is a positive probability that every element in the state space eventually gets mapped to any set of positive $\nu$ measure. It is $\phi$-recurrent if, with probability 1, it visits every set of positive $\phi$-measure in a finite number of steps.
The Krylov-Bogolubov Theorem [3] tells us that a stochastic dynamical system obtained from a continuous function on a compact subset of ${\mathbb{R}}^{d}$ with independent parametric noise always has an invariant distribution:
Theorem 1 (Krylov-Bogolubov).
A Feller continuous, stationary Markov process on a compact subset of ${\mathbb{R}}^{d}$ has an invariant distribution.
The theorem applies to more general settings, but this more specific version is all we need here. Irreducibility can then be used to show that this invariant distribution is unique using the following theorem:
@@
[[14], Theorem C9.3 in Chapter 2] If a Markov process is irreducible with respect to some $\sigma$-finite measure $\nu$ on $S$ and has an invariant probability $\mu^{*}$, then the process is positive Harris recurrent, $\mu^{*}$ is the unique invariant probability and, for any probability measure $\mu$ on S:
@@
Thus, if a Markov process is irreducible and has some invariant measure, we know that any measure converges to a unique invariant measure, $\mu^{*}$, under the operator, $P$ “in the sense of Césaro”, i.e. the average of the sequence of measures converges to $\mu^{*}$. We also get positive Harris recurrence from this result.
@@
A Markov process is said to be strongly aperiodic with respect to a $\sigma$-finite probability measure $\phi$ on $S$ if there is a set $A_{0}\subset S$ with $\phi(A_{0})=1$ and an $\alpha>0$ such that one has
@@
Suppose $\{X_{n}\}$ is a positive Harris recurrent and aperiodic Markov process with stationary distribution $\pi$. Then for any initial distribution, $\mu$, of $X_{0}$:
@@
We note that the measures used in the definitions of irreducibility and aperiodicity could be different. (This fact will be used later in the proof of Theorem 5.) To see that this is the case, consider the following. Aperiodicity with respect to the measure $\nu$ implies that there is an $\alpha>0$ such that
$p(x,B)\geq\alpha\nu(B)\,\Rightarrow\,p(x,B)>0\;\hbox{ if }\nu(B)>0$
Thus, the process is $\nu$-irreducible. Therefore, in the statement of Theorem 3, we do not need to specify the measures used to show either irreducibility or aperiodicity.
@@
where $\lambda$ is distributed according to a Bernoulli distribution, and takes two possible values, $\alpha$ and $\beta$ (where $\alpha<\beta$). Thus, the process switches randomly between two quadratic maps. Bhattacharya and Rao demonstrate that if $0\leq\alpha<\beta\leq 1$, then $\delta_{0}$, the Dirac measure supported on 0, is the unique invariant probability on [0,1]. If $1<\alpha<\beta\leq 2$ then there is a unique invariant probability on (0,1), and this probability is nonatomic. If $2<\alpha<\beta<1+\sqrt{5}$ and $\alpha\in[8/\beta(4-\beta),\beta)$, then there exists a unique invariant probability on (0,1) which is nonatomic and has its support contained in $[1/2,(1+\sqrt{5})/4]$.
@@
Recall that the deterministic logistic family: $f(x,\lambda)=f_{\lambda}(x)=\lambda x(1-x)$ undergoes a period-doubling cascade as $\lambda$ increases from 1 to $\lambda_{2\infty}\approx$ 3.56995. Note that here iteration occurs only with respect to $x$; that is $f^{(m)}(x,\lambda)=f(f^{m-1}(x,\lambda),\lambda)$, where $\lambda$ remains fixed. For some higher values of $\lambda$, the family exhibits stable periodic orbits of odd period, following the Sarkowski sequence, culminating with a stable period-3 orbit when $\lambda=1+\sqrt{8}$. Interspersed between the values of $\lambda$ with stable periodic orbits are a set of large measure of parameter values for which there are no stable periodic orbits, and almost all initial conditions give rise to orbits with no finite period. Thus, $\lambda_{2\infty}$ marks the “onset of chaos”. As $\lambda$ increases beyond $1+\sqrt{8}$, the family of maps again experiences a period-doubling cascade, with stable orbits of period 6, 12, 24, 48, …, culminating at $\lambda_{3\infty}\approx 3.8495$.
When $\lambda$ is greater than this value, the maps exhibit chaotic behavior on a Cantor set of parameter values, with three “chaotic bands” of orbits merging into one chaotic band at $\lambda\approx 3.857$. In what follows, we are interested in randomly perturbing the values of $\lambda$ in this chaotic regime. To make sure we are well within the parameter range for which there is generically a single band chaotic attractor, we restrict $\lambda\in(3.87,4)$. The logistic family is well-studied in the literature, and more details on its dynamics can be found in, for example, [17]
In [18], Athreya and Dai consider the stochastic logistic map and demonstrate that the Markov process $\{X_{n}\}$ is Feller continuous, which guarantees the existence of an invariant distribution (but not its stability) by Theorem 1. They also give conditions for convergence of the distribution of the $X_{i}$ in the sense of Césaro which differ somewhat from the hypotheses of Theorem 2. Here we choose to follow Bhattacharya and Majumbdar ([14]) and Nummelin ([15]) by showing the irreducibility and aperiodicity of the Markov process in the chaotic regime. This, along with the existence of an invariant measure, gives the asymptotic stability of the invariant measure by Theorem 3.
Lemma 1 and its generalization Lemma 2 build toward Theorem 4, which shows that the Markov process is irreducible. Theorem 5 shows that the process is strongly aperiodic and thus, since it has an invariant measure by Theorem 1, the process has a unique asymptotically stable invariant measure.
@@
Let $\lambda\sim\mathcal{U}(3.87,4)$, $\alpha$ a fixed element in $(3.87,4)$, and $x^{*}\in(0,1)$ such that $x^{*}=\alpha x^{*}(1-x^{*})$. (Equivalently, $x^{*}=1-\frac{1}{\alpha}$.) Then there exists $\eta>0$ such that for all $x\in I=(x^{*}-\eta,x^{*}+\eta)$ and $A\subset I$ with $m(A)>0$, then
$p(x,A)=\hbox{Prob}\{X_{1}\in A\,|\,X_{0}=x\}>0.$
where $m$ is Lebesgue measure.
That is, there is a neighborhood $I$ around the fixed point for which every $x\in I$ has a positive probability of visiting each subset of $I$ of positive Lebesgue measure.
@@
Fix $\alpha\in(3.87,4)$, and let $x^{*}\in(0,1)$ be such that $\alpha x^{*}(1-x^{*})=x^{*}$.
Notice that for any $\eta>0$ and $A\subset I\equiv(x^{*}-\eta,x^{*}+\eta)$ with $m(A)>0$
[…]
@@
[…]
Since $h(x)=x(1-x)$ is continuous at $x^{*}$ and $\alpha h(x^{*})=x^{*}$, it follows that there exists $\eta^{*}>0$ such that $a(\eta^{*})<4$ and $b(\eta^{*})>3.87$.   Let $A_{\eta^{*}}\subset I_{\eta^{*}}=(x^{*}-\eta^{*},x^{*}+\eta^{*})$ be a set of positive Lebesgue measure associated with this particular value of $\eta$. Then, for any $x\in S$, ${\frac{A_{\eta^{*}}}{x(1-x)}}\subset(3.87,4)$, and since $A_{\eta^{*}}$ has positive measure, so does $\frac{A_{\eta^{*}}}{x(1-x)}$, and therefore $p(x,A_{\eta^{*}})>0$.
Remark 1.
Note that while we take the distribution of $\lambda$ on $(3.87,4)$ to be uniform, the same result will hold if we replace the uniform distribution with a distribution that has non-vanishing absolutely continuous part on this interval.
The next lemma generalizes Lemma 1 to the case in which the map has a stable periodic orbit of any period.
Lemma 2.
As before, let $\lambda_{i}\sim\mathcal{U}(3.87,4)$ for all $i\geq 1$. Suppose $\alpha$ is a parameter value in $(3.87,4)$ with an $M$-periodic point, $x^{*}$, so that $f^{(M)}(x^{*},\alpha)=x^{*}$. Then there is an $\eta>0$ such that, for any $x\in I=(x^{*}-\eta,x^{*}+\eta)$ and $A\subset I$ with $m(A)>0$, we have
$p^{M}(x,A)=Prob(X_{M}\in A|X_{0}=x)>0$
Proof.
Since $43.87<\alpha<$ 4 we can choose an $\eta_{1}$ small enough so that
$43.87<\left({\frac{x^{*}-\eta_{1}}{x^{*}+\alpha\eta_{1}}}\right)\alpha<\left({\frac{x^{*}+\eta_{1}}{x^{*}-\alpha\eta_{1}}}\right)\alpha<$ 4
because
$\lim_{\eta_{1}\rightarrow 0}\frac{x^{*}-\eta_{1}}{x^{*}+\alpha\eta_{1}}=\lim_{\eta_{1}\rightarrow 0}\frac{x^{*}+\eta_{1}}{x^{*}-\alpha\eta_{1}}=1$
Denote points in the periodic orbit of $x^{*}$ as $x^{*}_{j}=f^{(j)}(x^{*},\alpha)$. Since $x^{*}$ has period $M$, $f(x^{*}_{M-1},\alpha)=\alpha x^{*}_{M-1}(1-x^{*}_{M-1})=x^{*}$. Since $f(\cdot,\cdot)$ is continuous in both variables:
  1. 1.
    there is an $\eta_{2}>0$ such that, if $|y-x^{*}_{M-1}|<\eta_{2}$ then
    $|f(y,\alpha)-x^{*}|<\eta_{1}\alpha\Leftrightarrow\left|y(1-y)-\frac{x^{*}}{\alpha}\right|<\eta_{1}\,;$
  2. 2.
    given $\eta_{2}$ as in the previous statement, $\exists\,\eta_{3}$ so that $|x-x^{*}|<\eta_{3}$ and $|\lambda_{i}-\alpha|<\eta_{3}$ for $1\leq i<M\Rightarrow|X_{M-1}-x^{*}_{M-1}|<\eta_{2}$ where $X_{j+1}=\lambda_{j+1}X_{j}(1-X_{j})$ and $X_{0}=x$.
In other words, if the random logistic map starts close to $x^{*}$, and if all of the parameter values are close to $\alpha$, then the random orbit returns close to $x^{*}$ after $M$ iterates.
Let $0<\eta<\min\{\eta_{1},\eta_{3}\}$. Then, if $|\lambda_{i}-\alpha|<\eta$ for $1\leq i<M$, and $A\subset I=(p-\eta,p+\eta)$, with $X_{0}\in I$, we have:
$\frac{A}{X_{M-1}(1-X_{M-1})}\subset\left({\frac{p-\eta_{1}}{\frac{p}{\alpha}+\eta_{1}},\frac{p+\eta_{1}}{\frac{p}{\alpha}-\eta_{1}}}\right)\subset(3.87,4)$
by our original assumption on $\eta_{1}$. Since $A$ has positive measure, so does ${\frac{A}{X_{M-1}(1-X_{M-1})}}$. To help clarify the notation, let $C$ be the event that $\lambda_{i}\in(\alpha-\eta,\alpha+\eta)$ for $1\leq i<M$. What we have shown is that
$Prob(X_{M}\in A|X_{M-1},C)=Prob\left({\lambda_{M}\in\frac{A}{X_{M-1}(1-X_{M-1})}|X_{M-1},C}\right)>0.$ (3)
Assume that $X_{0}=x\in I$, and recall that $\lambda_{M}$ is independent of $X_{0},X_{1},\dots X_{M-1}$. Then by the law of total probability:
$Prob(X_{M}\in A)\geq Prob((X_{M}\in A)\cap C)=E_{x}[Prob((X_{M}\in A)\cap C)|(X_{M-1}=x)]$ (4)
where $E_{x}$ is the expected value with respect to the distribution of $X_{M-1}$. Focusing on what’s inside the expected value:
$Prob\left({(X_{M}\in A)\cap C|X_{M-1}}\right)=Prob(X_{M}\in A|X_{M-1},C)Prob(C)$
Since the $\lambda$’s are independent of $X_{i}$ and $Prob(|\lambda_{i}-\alpha|<\eta)>0$, we have by Equation (3)
$Prob\left({(X_{M}\in A)\cap C|X_{M-1}}\right)=Prob(X_{M}\in A|X_{M-1},C)\left[{Prob(|\lambda_{i}-\alpha|<\eta)}\right]^{M-1}>0$
The expected value of a strictly positive random variable is positive, so we conclude, using Equation (4) that, if $X_{0}\in I$ then $Prob(X_{M}\in A)>0$, as claimed. ∎
Theorem 4.
As above, let $S=[0,1]$.
$X_{n+1}=\lambda_{n+1}X_{n}(1-X_{n}),\,\,\,n\geq 0$
with $\{\lambda_{n}\}_{n\geq 1}$ being i.i.d., $\lambda_{i}\sim\mathcal{U}(3.87,4)$, and $X_{0}$ an independent $S$-valued r.v. Then $\{X_{n}\}$ is irreducible with respect to the Lebesgue measure $m$ on $S$ conditioned on $(x^{*}-\eta,x^{*}+\eta)$.
Proof.
By a result of Graczyk and Swiatek [19], the set of all $\lambda$ in $(0,4)$ such that $f(\cdot,\lambda)$ has a stable periodic orbit is dense. Thus, there is an $\alpha\in(3.87,4)$ and $10<x^{*}<$ 1 such that $x^{*}$ is a stable periodic point for $f(x,\alpha)=\alpha x(1-x)$.
Let $K=\{x\in(0,1)\,\,|\,\,\omega_{\alpha}(x)=\gamma(x^{*})\}$, where $\omega_{\alpha}$ is the limit set of the orbit of $x$ under $f(\cdot,\alpha)$ and $\gamma(x^{*})$ is the orbit of the stable periodic point $x^{*}$. In [20] (Chapter V, Theorem 1.3) de Melo and van Strien have shown that $m(K)=1$ for unimodal maps, where $m$ is Lebesgue measure. Thus let $I=(x^{*}-\eta,x^{*}+\eta)$ as given by Lemma 2. Then, for all $x\in K$, $f^{(n)}(x,\alpha)\in I$ for some large $n$. Thus, by continuity of $f$,
$Prob(X_{n}\in I\text{ for some }n\geq 1|X_{0}=x)>0$ (5)
for all $x\in K$.
We will next show that for any $x\in(0,1)$, there is a positive probability that $x$ is mapped to $K$ in one iterate. In fact, we will show that $p(x,K)=Prob(X_{1}\in K|X_{0}=x)=1$. Fix $x\in(0,1)$ and let $K^{c}$ be the complement of $K$ in $[0,1]$. Since $\lambda$ is uniformly distributed on $(3.87,4)$ and $m(K^{c})=0$ we have
$p(x,K)$ $=$ $Prob(\lambda\,|\,\lambda x(1-x)\in K)$
$=$ $Prob\left({\lambda\in\frac{K}{x(1-x)}}\right)$
$=$ $Prob\left({(3.87,4)\cap\frac{K}{x(1-x)}}\right)$
$=$ $1-Prob\left({(3.87,4)\cap\frac{K^{c}}{x(1-x)}}\right)$
$\geq$ $1\displaystyle 1-Prob\left({\frac{K^{c}}{x(1-x)}}\right)=$ 1
Therefore, $p(x,K)>0$, establishing irreducibility. ∎
Using Theorems 1 and 2, we now have that, since the process is Feller continuous by [18], the process $\{X_{i}\}$ has a unique invariant measure that is stable in the sense of Césaro. We conclude this section by demonstrating that the process is strongly aperiodic, and thus the invariant distribution is asymptotically stable with respect to the total variation norm.
Theorem 5.
Let $S=[0,1]$.
$X_{n+1}=\lambda_{n+1}X_{n}(1-X_{n}),\,\,\,n\geq 0$
with $\{\lambda_{n}\}_{n\geq 1}$ being i.i.d., $\lambda_{i}\sim\mathcal{U}(3.87,4)$, and $X_{0}$ an independent $S$-valued r.v. Then $\{X_{n}\}$ is strongly aperiodic, and thus has an asymptotically stable invariant distribution.
Proof.
Let $\lambda_{min}=3.87$, and let $A_{0}=(1-\frac{1}{\lambda_{min}},1-\frac{1}{4})$ (that is, $A_{0}$ is the set of fixed points for the deterministic logistic maps corresponding to the possible parameter $\lambda$ values). Let $L_{1}$ be the length of $A_{0}$ (that is, $L_{1}=\frac{1}{\lambda_{min}}-\frac{1}{4}$ ) and let $L_{2}$ be the length of interval $(\lambda_{min},4)$ (that is, $L_{2}=4-\lambda_{\min}$). Let $\nu$ be normalized Lebesgue measure on $A_{0}$. Then $A_{0}$, by definition, has measure 1, and for any measurable set $A$ :
[…]
Fix $x\in A_{0}$ and $A\subset S$. We want to show that $p(x,A)\geq c\,\nu(A)$ for some positive constant, $c$ (Definition 5). Since $\nu(A)=\nu(A\cap A_{0})$ it suffices to demonstrate the above result for sets $A\subset A_{0}$.
Notice for any $A\subset A_{0}$ and with $\nu(A)>0$, and $x\in A_{0}$,
[…]
and
[…]
@@
[…]
$=$ $\frac{1}{L_{2}}\cdot\frac{1}{x(1-x)}\cdot L_{1}\nu(A)$
$\geq$ $4\cdot\frac{L_{1}}{L_{2}}\nu(A),$
since $x(1-x)\leq\frac{1}{4}$ for all $x\in[0,1]$. Thus, letting $\alpha=4\cdot\frac{L_{1}}{L_{2}}$ gives us the desired result. Therefore, the process $\{X_{n}\}$ is strongly aperiodic, and has an asymptotically stable invariant distribution. ∎

### v4

@@ 2 Background and Definitions
We consider the logistic Markov process
@@
    The Markov process is $\nu$-recurrent if, for each $x\in S$ and $A\in\mathcal{B}(S)$ such that $\nu(A)>0$, we have
@@
    If the Markov process is both $\nu$-irreducible and $\nu$-recurrent, then it is Harris ($\nu$-) recurrent.
@@
    for some $\nu$-recurrent set $A_{0}$, then the process is said to be positive Harris ($\nu$) recurrent.
In other words: a Markov process is $\nu$-irreducible if there is a positive probability that every element in the state space eventually gets mapped to any set of positive $\nu$ measure. It is $\nu$-recurrent if, with probability 1, it visits every set of positive $\nu$-measure in a finite number of steps.
Definition 3.
The set $A_{0}$ is $\nu$-recurrent if $\nu(A_{0})>0$ and it is recurrent, i.e. $P(\eta_{A_{0}}<\infty|X_{0}=x)=1$ $\forall x\in{\mathcal{B}}(S)$.
Irreducibility can then be used to show that the invariant distribution is unique using the following theorem:
@@
[[14], Theorem C9.3 in Chapter 2] If a Markov process is $\nu$-irreducible with respect to some $\sigma$-finite measure $\nu$ on $S$ and has an invariant probability $\mu^{*}$, then the process is positive Harris ($\mu^{*}$-) recurrent, $\mu^{*}$ is the unique invariant probability and, for any probability measure $\mu$ on S:
@@
Thus, if a Markov process is irreducible and has some invariant measure, we know that any measure converges to a unique invariant measure, $\mu^{*}$, under the operator, $P$ “in the sense of Césaro”, i.e. the average of the sequence of measures converges to $\mu^{*}$. Note that a result from [15] (Proposition 10.1.2) states that the invariant measure $\mu^{*}$ dominates the irreducible measure $\nu$ (that is, every set of positive $\nu$ measure also has positive $\mu^{*}$ measure), so Theorem 1 also implies positive Harris recurrence with respect to $\nu$.
@@
A Markov process is said to be strongly aperiodic with respect to a $\sigma$-finite probability measure $\nu$ on $S$ if there is a set $A_{0}\subset S$ with $\nu(A_{0})=1$ and an $c>0$ such that one has
@@
Suppose $\{X_{n}\}$ is a positive Harris recurrent and strongly aperiodic Markov process with respect to some nonzero $\sigma$-finite measure $\nu$, with stationary distribution $\pi$. Then for any initial distribution, $\mu$, of $X_{0}$:
@@
where $\lambda$ is distributed according to a Bernoulli distribution, and takes two possible values, $\alpha$ and $\beta$ (where $\alpha<\beta$). Thus, the process switches randomly between two quadratic maps. Bhattacharya and Rao demonstrate that if $0\leq\alpha<\beta\leq 1$, then $\delta_{0}$, the Dirac measure supported on 0, is the unique invariant probability on [0,1]. If $1<\alpha<\beta\leq 2$ then there is a unique invariant probability on (0,1), and this probability is nonatomic. If $2<\alpha<\beta<1+\sqrt{5}$ and $\alpha\in[8/\beta(4-\beta),\beta)$, then there exists a unique invariant probability on (0,1) which is nonatomic and has its support contained in $1/2,(1+\sqrt{5})/4]$.
@@
Recall that the deterministic logistic family: $f(x,\lambda)=f_{\lambda}(x)=\lambda x(1-x)$ undergoes a period-doubling cascade as $\lambda$ increases from 1 to $\lambda_{2^{\infty}}\approx$ 3.56995. Note that here iteration occurs only with respect to $x$; that is $f^{(m)}(x,\lambda)=f(f^{m-1}(x,\lambda),\lambda)$, where $\lambda$ remains fixed. For some higher values of $\lambda$, the family exhibits stable periodic orbits of odd period, following the Sarkowski sequence, culminating with a stable period-3 orbit when $\lambda=1+\sqrt{8}$. Interspersed between the values of $\lambda$ with stable periodic orbits are a set of large measure of parameter values for which there are no stable periodic orbits, and almost all initial conditions give rise to orbits with no finite period. Thus, $\lambda_{2^{\infty}}$ marks the “onset of chaos”. As $\lambda$ increases beyond $1+\sqrt{8}$, the family of maps again experiences a period-doubling cascade, with stable orbits of period 6, 12, 24, 48, …, culminating at $\lambda_{3^{\infty}}\approx 3.8495$.
When $\lambda$ is greater than this value, the maps exhibit chaotic behavior on a Cantor set of parameter values, with three “chaotic bands” of orbits merging into one chaotic band at $\lambda\approx 3.857$. In what follows, we are interested in randomly perturbing the values of $\lambda$ in this chaotic regime. To make sure we are well within the parameter range for which there is generically a single band chaotic attractor, we restrict $\lambda\in(3.87,4)$. The logistic family is well-studied in the literature, and more details on its dynamics can be found in, for example, [18].
We consider the stochastic logistic map, i.e. the Markov process given by
$X_{n+1}=\lambda_{n+1}X_{n}(1-X_{n})$ (3)
where $\{\lambda\}_{i\geq 1}$ are i.i.d. random variables distributed uniformly on (3.87,4). The initial state $X_{0}$ is a random variable taking values in (0,1), with distribution independent of the $\lambda$’s. In ([19] Theorem 2), Athreya and Dai demonstrate that the Markov process $\{X_{n}\}$ given by Equation (3) has an invariant distribution.
We structure this section as follows: we first show the irreducibility of the Markov process in the chaotic regime using Lemmas 1 and 2, which together with the existence of the invariant measure gives positive Harris recurrence, by Theorem 1. This, along with strong aperiodicity, gives the asymptotic stability of the invariant measure by Theorem 2.
The following lemma uses the existence of a dense orbit and continuity of the iterated functions to show that every $x\in(0,1)$ has a positive probability of entering any open set in finite time. This lemma will be used to prove the irreducibility of $\{X_{n}\}$.
@@
Let $A$ be a nonempty open subset of $(0,1).$ For all $x\in(0,1)$ there is an $M$ such that $p^{M}(x,A)=Prob(X_{M}\in A|X_{0}=x)>0$.
@@
It is known [20] that there is a value $\alpha\in(3.87,4)$ such that $f(x,\alpha)=\alpha x(1-x)$ exhibits topological transitivity; that is $f(x,\alpha)$ contains an orbit that is dense in $0,\alpha/4)$. Call this dense orbit $\Gamma.$ Let $x\in(0,1)$ Note then that, for all $x\in(0,1)$, the set $\{\lambda x(1-x):3.87<\lambda<4\}$ is an open interval intersecting $(0,\alpha/4)$, and therefore must contain a point on $\Gamma.$ Therefore, there exists $\beta\in(3.87,4)$ such that $y=f(x,\beta)$ is on $\Gamma$.
Note then, that since the orbit of $y$ is dense for parameter $\alpha$, there exists $M$ such that $f^{M}(y,\alpha)\in A$. Then, $f^{M}(f(x,\beta),\alpha)\in A$. Note further that $f(\ldots,f(x,\lambda_{1}),\cdots,\lambda_{M+1})$ is a polynomial in $M+2$ variables and is thus continuous everywhere. Therefore, since $f^{M}(f(x,\beta),\alpha)\in A$, and $A$ is an open set, there is a set of positive normalized Lebesgue measure $U$ in $(3.87,4)^{M+1}$ containing $(\beta,\underbrace{\alpha,\;\cdots\;,\alpha}_{\hbox{$M$ times}})$ such that for all $(\lambda_{1},\cdots,\lambda_{M+1})\in U$, $f(\ldots,f(x,\lambda_{1}),\cdots,\lambda_{M+1})\in A$. This means that
$Prob(X_{M+1}\in A|X_{0}=x)>0.$
The following lemma is used later to show strong aperiodicity as well as irreducibility.
Lemma 2.
Let $S=(0,1)$.
$X_{n+1}=\lambda_{n+1}X_{n}(1-X_{n}),\,\,\,n\geq 0$
with $\{\lambda_{n}\}_{n\geq 1}$ being i.i.d., $\lambda_{i}\sim\mathcal{U}(3.87,4)$, and $X_{0}$ an independent $S$-valued r.v. Let $\lambda_{min}=3.87$, and let $A_{0}=(1-\frac{1}{\lambda_{min}},1-\frac{1}{4})$ (that is, $A_{0}$ is the set of fixed points for the deterministic logistic maps corresponding to the possible parameter $\lambda$ values). Let $\phi$ be normalized Lebesgue measure on $A_{0}$. Then there exists $c>0$ such that for all $x\in A_{0}$, and $A$ a measurable subset of $(0,1)$, we have $p(x,A)\geq c\phi(A).$
Proof.
Let $\lambda_{min}=3.87$, and let $A_{0}=(1-\frac{1}{\lambda_{min}},1-\frac{1}{4})$ (that is, $A_{0}$ is the set of fixed points for the deterministic logistic maps corresponding to the possible parameter $\lambda$ values). Let $L_{1}$ be the length of $A_{0}$ (that is, $L_{1}=\frac{1}{\lambda_{min}}-\frac{1}{4}$ ) and let $L_{2}$ be the length of interval $(\lambda_{min},4)$ (that is, $L_{2}=4-\lambda_{\min}$). Let $\phi$ be normalized Lebesgue measure on $A_{0}$. Then $A_{0}$, by definition, has measure 1, and for any measurable set $A$ :
$\phi(A)=\frac{1}{L_{1}}\int_{y\in(A\cap A_{0})}dy$
Fix $x\in A_{0}$ and $A\subset S$. We want to show that $p(x,A)\geq c\,\phi(A)$ for some positive constant, $c$ (Definition 4). Since $\phi(A)=\phi(A\cap A_{0})$ it suffices to demonstrate the above result for sets $A\subset A_{0}$.
Notice for any $A\subset A_{0}$ and with $\phi(A)>0$, and $x\in A_{0}$,
$\underset{x}{\sup}\left\{\frac{A}{x(1-x)}\right\}<\frac{1-\frac{1}{4}}{\frac{1}{4}\left(1-\frac{1}{4}\right)}=4$
@@
$\underset{x}{\inf}\left\{\frac{A}{x(1-x)}\right\}>\frac{1-\frac{1}{\lambda_{\min}}}{\frac{1}{\lambda_{\min}}\left(1-\frac{1}{\lambda_{\min}}\right)}=\lambda_{\min}=3.87$
@@
$=$ $\displaystyle\frac{1}{L_{2}}\int_{\frac{A}{x(1-x)}}d\lambda$
$=$ $\frac{1}{L_{2}}\int_{A}\frac{1}{x(1-x)}dy$
$=$ $\frac{1}{L_{2}}\frac{1}{x(1-x)}\int_{A}dy$
$=$ $\frac{1}{L_{2}}\cdot\frac{1}{x(1-x)}\cdot L_{1}\phi(A)$
$\geq$ $4\cdot\frac{L_{1}}{L_{2}}\phi(A)$
$=$ $\frac{1}{\lambda_{\min}}\phi(A)\,$,
since $x(1-x)\leq\frac{1}{4}$ for all $x\in[0,1]$. Thus, letting $c=\displaystyle{\frac{1}{\lambda_{\min}}}$ gives us the desired result. ∎
Theorem 3.
Let $S=(0,1)$, and set $A_{0}$ be defined as above. As before, define the Markov process on $S$:
$X_{n+1}=\lambda_{n+1}X_{n}(1-X_{n}),\,\,\,n\geq 0$
with $\{\lambda_{n}\}_{n\geq 1}$ being i.i.d., $\lambda_{i}\sim\mathcal{U}(3.87,4)$, and $X_{0}$ an independent $S$-valued r.v. Then $\{X_{n}\}_{n=0}^{\infty}$ is irreducible with respect to normalized Lebesgue measure on $A_{0}$.
Proof.
Let $x\in(0,1)$, and let $B$ be a subset of positive measure. We’d like to show that there is some $n\geq 1$ such that $p^{n}(x,B)=P(X_{n}\in B|X_{0}=x)>0.$ Note that
$P(X_{n}\in B|X_{0}=x)=P(X_{n}\in B|X_{n-1}\in A_{0})*P(X_{n-1}\in A_{0}|X_{0}=x).$
By Lemma 2, $P(X_{n}\in B|X_{n-1}\in A_{0})\geq c\,\phi(B)>0$, where $c=\frac{1}{\lambda_{min}}$. $P(X_{n-1}\in A_{0}|X_{0}=x)>0$ by Lemma 1, since $A_{0}$ is an open interval. Thus, $p^{n}(x,B)=P(X_{n}\in B|X_{0}=x)>0$ and $\{X_{n}\}_{n=0}^{\infty}$ is irreducible with respect to normalized Lebesgue measure on $A_{0}$, ∎
By the existence of the invariant measure by [19] and Theorem 1, the process is positive Harris recurrent with respect to $\phi$, and the process $\{X_{i}\}$ has a unique invariant measure that is stable in the sense of Césaro. We conclude this section by demonstrating that the process is strongly aperiodic, and thus the invariant distribution is asymptotically stable with respect to the total variation norm.
Theorem 4.
Let $S=(0,1)$.
$X_{n+1}=\lambda_{n+1}X_{n}(1-X_{n}),\,\,\,n\geq 0$
with $\{\lambda_{n}\}_{n\geq 1}$ being i.i.d., $\lambda_{i}\sim\mathcal{U}(3.87,4)$, and $X_{0}$ an independent $S$-valued r.v. Then $\{X_{n}\}_{n=0}^{\infty}$ has an invariant distribution that is asymptotically stable with respect to the total variation norm.
Proof.
Because $\{X_{n}\}_{n=0}^{\infty}$ is positive Harris recurrent with respect to $\phi$, it is $\phi$-recurrent. This together with Lemma 2 gives that the process is strongly aperiodic with respect to $\phi$. By Theorem 2, this implies that $\{X_{n}\}$ has an invariant distribution that is asymptotically stable with respect to the total variation norm.